OpenStudy (anonymous):

Use the distance formula to find the equation of a parabola with focus (-2, 0) and directrix x = 8.

OpenStudy (amistre64):

umm ... whats preventing you from finishing it?

OpenStudy (amistre64):

for some point (x,y), the distance from the focus is equal to the perp distance to the directrix... (df)^2 = (-2-x)^2 + (0-y)^2 (dd)^2 = (8-x)^2 + (y-y)^2 when does (df)^2 = (dd)^2 ? (-2-x)^2 + y^2 = (8-x)^2